# Identifying and Mitigating API Misuse in Large Language Models

### v1

@@ Abstract
API misuse in code generated by large language models (LLMs) represents a serious emerging challenge in software development. While LLMs have demonstrated impressive code generation capabilities, their interactions with complex library APIs remain highly prone to errors, potentially leading to software failures and security vulnerabilities. This paper presents the first comprehensive study of API misuse patterns in LLM-generated code, analyzing both method selection and parameter usage across Python and Java. Through extensive manual annotation of 3,892 method-level and 2,560 parameter-level misuses, we develop a novel taxonomy of four distinct API misuse types specific to LLMs, which significantly differ from traditional human-centric misuse patterns. Our evaluation of two widely used LLMs, StarCoder-7B (open-source) and Copilot (closed-source), reveals significant challenges in API usage, particularly in areas of hallucination and intent misalignment. We propose Dr.Fix, a novel LLM-based automatic program repair approach for API misuse based on the aforementioned taxonomy. Our method substantially improves repair accuracy for real-world API misuse, demonstrated by increases of up to 38.4 points in BLEU scores and 40 percentage points in exact match rates across different models and programming languages. This work provides crucial insights into the limitations of current LLMs in API usage and presents an effective solution for the automated repair of API misuse in LLM-generated code.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly enhanced the sophistication of automatic code generation, as these models are trained on vast datasets of source code [1, 2, 3, 4]. However, according to previous work, LLMs may still generate suboptimal or incorrect API usage in code generation despite being exposed to extensive code during training [5]. Library APIs have been pivotal in software development, enabling complex functionalities across various domains, including network communication, data visualization, and scientific computation [6]. Moreover, as libraries evolve by introducing new features, deprecating old ones, or changing usage patterns, LLMs may struggle to adapt to these changes [7, 8]. With over 590,000 external libraries [^1] https://pypi.org/ in languages like Python, using the right API correctly can be a considerable challenge, even for experienced developers.
@@
Despite these advancements, existing studies on API misuse in LLM-generated code have focused primarily on controlled, synthetic scenarios. For example, [9] investigated the quality of code generated by several LLMs using StackOverflow-style programming questions, finding that GPT-4 generated code with API misuses in 62% of cases when tested on 18 widely used Java APIs. Similarly, [10] studied API misuse in LLMs by examining security-related programming questions in Java, revealing that LLMs misused APIs in approximately half of the tasks. A more recent error analysis of 1,000 samples from existing benchmarks identified three primary error types, “AttributeError”, “TypeError”, and “ValueError”, commonly associated with API misuse in LLM-generated code. While these studies have demonstrated that specific LLMs can generate code with API misuse on limited libraries, APIs, and programming languages, their findings may not generalize to real-world programming contexts due to their reliance on human-curated questions and scenarios.
@@
A more common scenario in real-world software development involves LLMs integrated within IDE environments, which assist developers in continuing or completing existing code [11]. LLMs are often prompted with an incomplete code snippet and asked to suggest an appropriate API call. In this context, LLMs face distinct challenges: They may erroneously hallucinate functions from other libraries, suggest irrelevant or incorrect APIs, or specify invalid parameters, even when familiar with the target library. These context-dependent completion tasks present challenges different from those of generating standalone code snippets in response to specific queries. While previous research has studied API misuse of LLMs on task-oriented synthetic datasets [9, 10], there remains a critical gap in understanding whether LLMs exhibit similar patterns of API misuse in real-world code completion. Specifically, it remains unclear whether LLMs make the same types of mistakes in API usage as human developers do. Moreover, investigating API misuse patterns across both open-source and closed-source models is essential, as open-source models often lag behind their closed-source counterparts in terms of performance. Current approaches to mitigating API misuse primarily focus on automatic detection using predefined misuse types and API documentation. However, these methods face significant limitations when applied to LLM-based code generation, which involves millions of APIs and varying contexts. Even when misuse is detected correctly, rule- and program-analysis-based automatic program repair (APR) techniques often fail to fix misused APIs due to their limited semantic understanding [12, 13]. This suggests that LLM-based APR approaches, which can leverage the semantic understanding of code, may be more suitable for addressing such issues.
@@
This work investigates how LLMs misuse APIs and examines their behavior in real-world code-generation scenarios. Specifically, we evaluate two LLMs: StarCoder-7B [3], an open-source LLM for code completion, and Copilot, a fine-tuned closed-source model based on GPT-4 [2]. Both models are widely used in IDE environments such as VSCode [^2] https://code.visualstudio.com/. While previous human-centric API misuse research often studies patches from fixed commits [12, 13], it is not feasible to treat LLM-generated code in the same way, as LLMs generate outputs in various formats. We sample code snippets from GitHub in Python and Java to systematically assess how these models misuse APIs. We conduct program analysis to localize the API positions and provide the code context before these APIs as input to the models. Unlike existing studies, we assess the correctness of API invocation in two key areas: method selection and parameter usage. The first aspect evaluates whether LLMs can suggest an appropriate API name based on the code context, while the second examines whether LLMs can predict reasonable API parameters. These aspects correspond to typical use cases in IDE environments’ code filling [14] and left-to-right code completion [15].
@@
To the best of our knowledge, we are the first to analyze complex API misuse in LLM-based code generation through extensive manual annotation of thousands of code snippets. Our systematic study reveals that LLMs struggle significantly with API misuse, particularly hallucination and intent misalignment, with distinct patterns occurring between Python and Java. While Copilot outperforms StarCoder, their shared error patterns indicate common challenges in API usage. To address these issues, we propose a novel LLM-based APR method that significantly improves repair accuracy. BLEU scores increase by up to 38.4 points, and exact match rates improve by up to 40 percentage points across different models and languages. The key contributions of this research include:
  • We propose the first taxonomy of API misuse in the era of LLMs, which covers four types, different from human-centric misuse. We identify (1) Intent misuse, where models select syntactically correct but functionally inappropriate APIs; (2) Hallucination misuse, where models generate non-existent API methods or parameters; (3) Missing item misuse, where models omit required API parameters or methods, and (4) Redundancy misuse, where models introduce unnecessary API calls or method chaining that may lead to inefficiencies or errors.
  • We manually annotated the LLM-generated outputs and provided a dataset that contains 3892 method-level misuses (1,884 in Java and 2,048 in Python) and 2560 parameter-level misuses (1,279 in Java and 1,281 in Python). For each misuse sample, we provide the category based on the taxonomy.
  • We design a new LLM-based approach, Detect-reason-Fix (Dr.Fix) to automatically repair the misuse cases introduced by LLMs, demonstrating its efficacy among strong LLM-based APR methods.
@@
The rest of the paper is organized as follows. We outline the background in Section 2. We elaborate on the details of the experimental setting for the case study and introduce the misuse taxonomy in Section 3. Based on the annotations on generating two LLMs on Python and Java, we conduct quantitative analysis and share several insights in Section 4. Then, we propose the Dr.Fix for the in-wild API misuse repair and evaluate its performance against LLM-based APR on our benchmark in Section 5.
@@
import torch.nn as nn
class GRUCell(nn.Module):
@@
# LLM is expected to complete the next line with an API call, such as nn.Linear
@@ 2.1 Motivation
Consider the example of LLM code generation shown in Figure 1. The LLM is prompted with an incomplete code snippet and asked to suggest an appropriate API call. The LLM typically makes several errors when suggesting code completion or insertion: the LLM generates incomplete method calls and/or with erroneous parameters; the LLM generates method calls to similar but unreleated API; the LLM generates extraneous method calls not needed in the target usage context; the LLM does not follow a required sequence of method calls to properly set up and use the target library due to its API’s complexity; and the LLM needs to invoke API calls to multiple libraries but fails to correctly use these different library APIs together. All of these errors can be very challenging to detect and fix, statically and dynamically.
@@ 2.2 API Misuse
APIs are essential in modern software systems, enabling developers to interact with third-party libraries and services. Correct API usage ensures that the functionality of a library is used as intended, while misuse can lead to errors, unexpected behavior, inefficiencies, or security vulnerabilities [16]. API misuse occurs when developers violate the intended usage constraints of an API, such as input types, function call sequences, or data handling practices. Unlike general programming errors (e.g., type mismatches, null pointer dereferencing) [17], which are language-specific, API misuse may not immediately result in failure but can cause subtle issues like incorrect outputs or degraded performance [18].
The causes of API misuse are multifaceted, including unclear documentation, lack of domain-specific knowledge, and breaking changes in evolving APIs [19]. The consequences can range from minor performance bottlenecks to severe failures, especially in high-risk domains like healthcare or finance [10]. Addressing API misuse requires better documentation, improved tooling, and clear guidelines to ensure developers follow best practices. Research typically focuses on external misuse of libraries in open-source codebases, often identified by manually inspecting bug-fixing commits [20, 21, 22].
In this study, we focus on the misuse of third-party library APIs arising from improper usage patterns or violations of API contracts rather than issues related to the underlying programming language or general software evolution. We adopt definitions from [23], with minor adaptations. An API is the public interface exposing software elements (e.g., classes and methods) to the outside world. An API element includes any constituent part of an API used in client code, such as methods and parameters. An API misuse is the violation of one or more API usage constraints, leading to misbehavior such as errors, crashes, incorrect output, or vulnerabilities. Proper understanding and adherence to API constraints are crucial for preventing these issues and ensuring the API functions as intended.
@@ 2.3 Mitigation of API Misuse
API usage can be categorized into two primary types: directly calling API methods or instantiating objects from API classes [24]. API misuses are defined as violations of the implicit or explicit usage constraints of APIs [16]. Recently, numerous API-misuse detectors have been developed [25, 26, 27]. Generally, there are three types of API misuse detection approaches: static detectors, which identify API misuses through static analysis of source or binary code [28]; dynamic detectors, which detect API misuses through dynamic analysis [29]; and hybrid approaches that combine mining techniques with static analysis [30]. Regardless of the detection technique, existing methods (1) require API specifications, considering the violation of these specifications as API misuse. However, these specifications are often incomplete and difficult to obtain [27]; or (2) do not require explicit API specifications, instead relying on the assumption that the majority usage pattern in a large-scale code corpus represents valid usage. An API is considered misused if it deviates from this majority pattern. The limitation is that this assumption may not always hold, especially for rarely used APIs where the majority pattern does not necessarily reflect valid usage [26]. Recently, there has been promising work exploring the use of LLMs to repair API misuse by fine-tuning paired commit data [13]. This approach outperforms previous automatic program repair methods, providing a more effective solution to mitigate API misuse in code generation.
3 Taxonomy and Qualitative Studies
This section introduces a new taxonomy for API misuse in LLMs, based on our qualitative studies. We begin by establishing the framework for code generation and then exploring various types of API misuse under the proposed taxonomy.
@@ 3.1 Research Questions
RQ1: How well do LLMs handle API invocation in code generation? We first wanted to investigate how well LLMs can correctly invoke APIs in code generation tasks, using widely-used automatic evaluation metrics.

### v2

@@ Abstract
API misuse in code generated by large language models (LLMs) presents a serious and growing challenge in software development. While LLMs demonstrate impressive code generation capabilities, their interactions with complex library APIs are often error-prone, potentially leading to software failures and vulnerabilities. In this paper, we conduct a large-scale study of API misuse patterns in LLM-generated code, analyzing both method selection and parameter usage across Python and Java, using three representative LLMs (StarCoder-7B, Qwen2.5-Coder-7B, and GitHub Copilot). Based on extensive manual annotation of 3,209 method-level and 3,492 parameter-level misuses, we identify and categorize four recurring misuse types by building on and refining prior API misuse taxonomies. Our evaluation of three widely used LLMs, StarCoder-7B, Qwen2.5-Coder-7B, and GitHub Copilot, reveals persistent challenges in API usage, particularly hallucination and intent misalignment. To address these issues, we propose Dr.Fix, an LLM-based automatic repair approach guided by our taxonomy. Dr.Fix improves repair accuracy compared to baseline prompting and existing repair methods, with gains of up to 38.4 BLEU and 40% in exact match on benchmark datasets. This work offers important insights into the current limitations of LLMs in API usage and provides insights into current limitations and points to directions for improving automated misuse repair in code generation systems.
Index Terms:
Code Generation, Large Language Model, API Misuse, Automatic Program Repair, Empirical Software Engineering
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly enhanced the sophistication of automatic code generation, as these models are trained on vast datasets of source code [touvron2023llama, achiam2023gpt, li2023starcoder, lozhkov2024starcoder, muennighoff2024octopack, zhauo-etal-2025-nlp]. However, according to previous work, LLMs may still generate suboptimal or incorrect API usage in code generation despite being exposed to extensive code during training [zhuo2024bigcodebench]. Library APIs have been pivotal in software development, enabling complex functionalities across various domains, including network communication, data visualization, and scientific computation [zhong2017empirical]. Moreover, as libraries evolve by introducing new features, deprecating old ones, or changing usage patterns, LLMs may struggle to adapt to these changes [islah2024gitchameleon, kuhar2024libevolutioneval]. With over 590,000 external libraries [^1] https://pypi.org/ in languages like Python, using the right API correctly can be a considerable challenge, even for experienced developers.
@@
Despite these advancements, existing studies on API misuse in LLM-generated code have focused primarily on controlled, synthetic scenarios. For example, [zhong2024can] investigated the quality of code generated by several LLMs using StackOverflow-style programming questions, finding that GPT-4 generated code with API misuses in 62% of cases when tested on 18 widely used Java APIs. Similarly, [mousavi2024investigation] studied API misuse in LLMs by examining security-related programming questions in Java, revealing that LLMs misused APIs in approximately half of the tasks. A more recent error analysis of 1,000 samples from existing benchmarks identified three primary error types, “AttributeError”, “TypeError”, and “ValueError”, commonly associated with API misuse in LLM-generated code. While these studies have demonstrated that specific LLMs can generate code with API misuse on limited libraries, APIs, and programming languages, their findings may not generalize to real-world programming contexts due to their reliance on human-curated questions and scenarios.
@@
A more common scenario in real-world software development involves LLMs integrated within IDE environments, which assist developers in continuing or completing existing code [xu2022ide]. LLMs are often prompted with an incomplete code snippet and asked to suggest an appropriate API call. In this context, LLMs face distinct challenges: They may erroneously hallucinate functions from other libraries, suggest irrelevant or incorrect APIs, or specify invalid parameters, even when familiar with the target library. These context-dependent completion tasks present challenges different from those of generating standalone code snippets in response to specific queries. While previous research has studied API misuse of LLMs on task-oriented synthetic datasets [zhong2024can, mousavi2024investigation], there remains a critical gap in understanding whether LLMs exhibit similar patterns of API misuse in real-world code completion. Specifically, it remains unclear whether LLMs make the same types of mistakes in API usage as human developers do. Moreover, investigating API misuse patterns across both open-source and closed-source models is essential, as open-source models often lag behind their closed-source counterparts in terms of performance. Current approaches to mitigating API misuse primarily focus on automatic detection using predefined misuse types and API documentation. However, these methods face significant limitations when applied to LLM-based code generation, which involves millions of APIs and varying contexts. Even when misuse is detected correctly, rule- and program-analysis-based automatic program repair (APR) techniques often fail to fix misused APIs due to their limited semantic understanding [kechagia2021evaluating, zhang2023evaluating]. This suggests that LLM-based APR approaches, which can leverage the semantic understanding of code, may be more suitable for addressing such issues.
@@
In this work, we investigate how LLMs misuse APIs and examine their behavior in real-world code-generation scenarios. Specifically, we evaluate three representative decoder-only LLMs used in IDE settings: StarCoder-7B [li2023starcoder] and Qwen2.5-Coder-7B, two open-source models for code completion, and Copilot, a closed-source model powered by GPT-4 [achiam2023gpt]. Both models are widely used in IDE environments such as VSCode [^2] https://code.visualstudio.com/ . While previous human-centric API misuse research often studies patches from fixed commits [kechagia2021evaluating, zhang2023evaluating], it is not feasible to treat LLM-generated code in the same way, as LLMs generate outputs in various formats. We sample code snippets from GitHub in Python and Java to systematically assess how these models misuse APIs. We conduct program analysis to localize the API positions and provide the code context before these APIs as input to the models. Unlike existing studies, we assess the correctness of API invocation in two key areas: method selection and parameter usage. The first aspect evaluates whether LLMs can suggest an appropriate API name based on the code context, while the second examines whether LLMs can predict reasonable API parameters. These aspects correspond to typical use cases in IDE environments’ code filling [friedincoder] and left-to-right code completion [radford2018improving].
Building on existing taxonomies of API misuse [wei2024demystifying, he2023python], we adapt and refine them to the context of LLM-generated code. Specifically, we retain violation types that can be reliably identified from snippet-level context, such as Missing and Redundant, and introduce two additional categories that capture LLM-specific failure modes, namely Intent misuse and Hallucination misuse, while de-emphasizing categories that depend on broader project or version information, such as Replacement and Outdated. Through extensive manual annotation of 3,209 method-level and 3,492 parameter-level cases, we identify four recurring misuse patterns in LLM-generated API code: (1) Intent misuse, where a model selects an API that is syntactically valid but semantically inappropriate for the intended task; (2) Hallucination misuse, where the model produces non-existent or fabricated API methods or parameters; (3) Missing item misuse, where required API methods or parameters are omitted; and (4) Redundancy misuse, where unnecessary API calls or arguments are introduced, often leading to inefficiency or potential errors.
@@
Our systematic study reveals that LLMs struggle significantly with API misuse, particularly hallucination and intent misalignment, with distinct patterns observed between Python and Java. While Copilot and Qwen2.5-Coder generally outperform StarCoder, their shared error patterns indicate common challenges in API usage across models. To address these issues, we propose Detect-reason-Fix (Dr.Fix), a taxonomy-guided LLM-based APR method. Compared with baseline prompting and prior repair methods, Dr.Fix achieves substantial improvements, with BLEU scores increasing by up to 38.4 points and exact match rates improving by up to 40 percentage points across different models and languages.
The key contributions of this research include:
  • An empirical study of API misuse in LLM-generated code completion, conducted on Python and Java using three representative decoder-only LLMs (StarCoder-7B, Qwen2.5-Coder-7B, and Copilot).
  • An adaptation of existing API misuse taxonomies to the LLM setting, refined through the manual annotation of 6,452 misuse cases across method and parameter levels.
  • A publicly available dataset and replication package to support transparency and reproducibility.
  • A taxonomy-guided LLM-based APR approach, Dr.Fix, evaluated against baseline prompting and prior work, showing significant improvements in repair accuracy.
@@
The rest of the paper is organized as follows. We outline the background in Section 2. We elaborate on the details of the experimental setting for the case study and introduce the misuse taxonomy in Section 3. Based on the annotations on generating two LLMs on Python and Java, we conduct quantitative analysis and share several insights. Then, we propose the Dr.Fix for in-the-wild API misuse repair and evaluate its performance against LLM-based APR on our benchmark in Section 5. To facilitate the future research of API misuse, we provide a replication package [llm_api_misuse_replication] and plan to release the code and dataset publicly on GitHub upon acceptance.
@@
import torch.nn as $n$
class GRUCell($n$.Module):
@@
# LLM is expected to complete the next line with an API call, such as $n$.Linear
@@ 2.1 Motivation
Consider the example of LLM code generation shown in Figure 1. When prompted with an incomplete code snippet, an LLM may suggest an inappropriate API call due to misunderstanding the usage context. Typical errors include: (1) incomplete method calls or those with erroneous parameters, (2) calls to similar but unrelated APIs, (3) extraneous method calls not needed in the current context, (4) incorrect sequencing of method calls for library setup, and (5) incorrect integration of APIs from multiple libraries.
With the rise of support for both code completion and code infilling in modern IDEs such as Cursor [^3] https://docs.cursor.com/en/tab/overview , it is important to examine how reliably LLMs can generate correct API usage under these two scenarios. Since API calls are the central building blocks of modern software development, even small deviations from correct usage can propagate into functional errors, security issues, or inefficiencies. Therefore, before presenting our evaluation setup, we first scope what we consider to be API misuse in the context of LLM-generated code and how it differs from general programming errors.
@@ 2.2 API Misuse
APIs are essential components of modern software systems, enabling developers to interact with third-party libraries and services. Correct usage of APIs ensures that functionality is accessed as intended, while misuse can result in unexpected behavior, degraded performance, or even security vulnerabilities [amann2018systematic]. API misuse typically occurs when developers violate usage constraints, such as required input types, parameter ordering, or call sequences. Unlike general syntax or language-level type errors (e.g., null dereferencing [bush2000static]), API misuses may not always lead to immediate failure but can still cause subtle, incorrect behavior [khatchadourian2020empirical]. The causes of API misuse are multifaceted, including incomplete documentation, lack of domain knowledge, and evolving API designs [lamothe2021systematic]. Consequences range from performance degradation to system failure, particularly in high-stakes domains like healthcare or finance [mousavi2024investigation]. Prior studies have extensively investigated human-written API misuses through bug-fixing commits and empirical analysis [schlichtig2022fum, ren2020demystify, galappaththi2024empirical, wei2024demystifying].
In this work, we adopt the usage constraint framework from Schlichtig et al. [schlichtig2022fum] and apply the taxonomy proposed by Wei et al. [wei2024demystifying] and He et al. [he2023python], which classifies API misuses based on violation types (e.g., Missing, Redundant) and API elements (e.g., Method, Parameter, Condition). Although originally developed for deep learning libraries, we generalize this taxonomy to broader third-party APIs. We follow a closed coding approach to categorize LLM-generated misuses under this schema. Following [amann2016mubench], we define an API misuse as an incorrect use of an API that violates its documented contract or commonly expected usage constraints at the level of a specific API element. For instance, passing a string where a list is required is classified as an API misuse, even if the code is syntactically valid in Python, since it semantically violates the API’s intended usage. In contrast, general language-level errors such as undefined variables, misspellings, or unrelated type mismatches are excluded unless they directly affect API invocation [amann2016mubench, wei2024demystifying, he2023python]. We adapt this taxonomy for the LLM setting and introduce two descriptive refinements, intent misuse and hallucination, as surface-level subtypes of existing misuse categories. Intent misuse reflects cases where the model selects a valid API element that is semantically incorrect given the task (e.g., using vs.abs instead of a vector magnitude function). Hallucination captures cases where the model generates nonexistent methods or parameters. While both map to known categories like Replacement or Missing–Method, they are particularly prevalent in LLM outputs and not emphasized in prior human-centric taxonomies. We include them to better reflect distinctive patterns in LLM-generated code without altering the underlying taxonomy.
@@ 2.3 Mitigation of API Misuse
API usage can be categorized into two primary types: directly calling API methods or instantiating objects from API classes [nguyen2010graph]. API misuses are defined as violations of the implicit or explicit usage constraints of APIs [amann2018systematic]. Recently, numerous API-misuse detectors have been developed [kechagia2019effective, kang2021active, li2021arbitrar]. Generally, there are three types of API misuse detection approaches: static detectors, which identify API misuses through static analysis of source or binary code [monperrus2013detecting]; dynamic detectors, which detect API misuses through dynamic analysis [pradel2012leveraging]; and hybrid approaches that combine mining techniques with static analysis [pradel2012statically]. Regardless of the detection technique, existing methods (1) require API specifications, considering the violation of these specifications as API misuse. However, these specifications are often incomplete and difficult to obtain [li2021arbitrar]; or (2) do not require explicit API specifications, instead relying on the assumption that the majority usage pattern in a large-scale code corpus represents valid usage. An API is considered misused if it deviates from this majority pattern. The limitation is that this assumption may not always hold, especially for rarely used APIs where the majority pattern does not necessarily reflect valid usage [kang2021active]. More recently, there has been promising work exploring the use of LLMs to repair API misuse by fine-tuning paired commit data [zhang2023evaluating]. Other approaches, such as [wei2024demystifying, kechagia2021evaluating], provide domain-specific or rule-based repair mechanisms. In this work, we compare our proposed method not only against prompting baselines, but also contextualize it with these prior repair techniques, highlighting the advantages and limitations of LLM-based approaches for API misuse repair.
3 Qualitative Studies
This section presents our qualitative analysis of API misuse in LLM-generated code. We do not propose a new taxonomy; rather, we apply a well-established classification framework from Schlichtig et al. [schlichtig2022fum], supplemented with categories from Wei et al. [wei2024demystifying] and He et al. [he2023python], to annotate API misuse cases. These taxonomies define violations in terms of location (method, parameter, condition) and type (missing, redundant, replacement, outdated). Our goal is to assess whether LLMs exhibit similar misuse patterns to human developers, and where they differ.
We use this framework to analyze outputs from LLMs in realistic code completion tasks, focusing on API correctness in two common generation scenarios.
@@ 3.1 Research Questions
RQ1: How well do LLMs handle API invocation in code generation? We first investigate how reliably LLMs can generate correct API usages in two modern IDE scenarios: code completion, where models predict the next line or token sequence, and code infilling, where models fill in masked elements such as method names or parameters. These objectives reflect real-world developer interactions, where models are often used to suggest API methods or complete parameter lists within partially written code. Our design builds on traditional studies of code completion [proksch2016evaluating, raychev2014code] as well as recent work on infilling techniques [friedincoder, li2023starcoder, roziere2023code], ensuring that our evaluation setup aligns with both established and emerging IDE practices.